Determine the cost of giving up cash discounts under each of the following terms of sale (assuming 365 day year) : Show all steps A. 2-10 Net 30 b.1-10 Net 30 c. 1-10 net 45 d. 3/10 net 90 e. 1/10 net 60 f. 3/10 net 30 g. 4/10 net 180
Answer #1

Cost of Trade Credit = [1 + Discount rate/(1-discount rate)]^[365/(Days taken-Discount period)] - 1

Case A:

Cost of Trade Credit = [1 + 0.02/(1-0.02)]^[365/(30-10)] - 1
Cost of Trade Credit = [1 + 0.02/0.98]^(365/20) - 1
Cost of Trade Credit = 1.02041^(18.25) - 1
Cost of Trade Credit = 1.4459 - 1
Cost of Trade Credit = 0.4459
Cost of Trade Credit = 44.59%

Case B:

Cost of Trade Credit = [1 + 0.01/(1-0.01)]^[365/(30-10)] - 1
Cost of Trade Credit = [1 + 0.01/0.99]^(365/20) - 1
Cost of Trade Credit = 1.01010^(18.25) - 1
Cost of Trade Credit = 1.2013 - 1
Cost of Trade Credit = 0.2013
Cost of Trade Credit = 20.13%

Case C:

Cost of Trade Credit = [1 + 0.01/(1-0.01)]^[365/(45-10)] - 1
Cost of Trade Credit = [1 + 0.01/0.99]^(365/35) - 1
Cost of Trade Credit = 1.01010^(10.4286) - 1
Cost of Trade Credit = 1.1105 - 1
Cost of Trade Credit = 0.1105
Cost of Trade Credit = 11.05%

Case D:

Cost of Trade Credit = [1 + 0.03/(1-0.03)]^[365/(90-10)] - 1
Cost of Trade Credit = [1 + 0.03/0.97]^(365/80) - 1
Cost of Trade Credit = 1.03093^(4.5625) - 1
Cost of Trade Credit = 1.1491 - 1
Cost of Trade Credit = 0.1491
Cost of Trade Credit = 14.91%

Case E:

Cost of Trade Credit = [1 + 0.01/(1-0.01)]^[365/(60-10)] - 1
Cost of Trade Credit = [1 + 0.01/0.99]^(365/50) - 1
Cost of Trade Credit = 1.01010^(7.3) - 1
Cost of Trade Credit = 1.0761 - 1
Cost of Trade Credit = 0.0761
Cost of Trade Credit = 7.61%

Case F:

Cost of Trade Credit = [1 + 0.03/(1-0.03)]^[365/(30-10)] - 1
Cost of Trade Credit = [1 + 0.03/0.97]^(365/20) - 1
Cost of Trade Credit = 1.03093^(18.25) - 1
Cost of Trade Credit = 1.7435 - 1
Cost of Trade Credit = 0.7435
Cost of Trade Credit = 74.35%

Case G:

Cost of Trade Credit = [1 + 0.04/(1-0.04)]^[365/(180-10)] - 1
Cost of Trade Credit = [1 + 0.04/0.96]^(365/170) - 1
Cost of Trade Credit = 1.04167^(2.1471) - 1
Cost of Trade Credit = 1.0916 - 1
Cost of Trade Credit = 0.0916
Cost of Trade Credit = 9.16%